OpenStudy (anonymous):

Represent the area of a triangle with a base of 2x + y and a height of 4z (A) 2z(2x+y) (B)2x(z+y) (C)2xyz (D)z(2x + y ) (E) 2z + 4xy

OpenStudy (anonymous):

area of triangle is o.5 *base * height

OpenStudy (anonymous):

subsitute the base given adn the height given in to the equation above

jagr2713 (jagr2713):

Welcome to OpenStudy @602, hope your experience here is Fantastic! If you haven't yet please take the time and look over our Code of conduct http://openstudy.com/code-of-conduct Feel free to message me anytime you need help.

OpenStudy (anonymous):

A to E are the choices

OpenStudy (anonymous):

\[ A = \frac 12 bh = \frac 12 (2x+y)(4z) = 2z(2x+y) \]

OpenStudy (anonymous):

Ty